OpenStudy (anonymous):

How many real zeroes does y = ( x + 8 )^3 + 9 have?

OpenStudy (sirm3d):

one only. the two other zeroes not real.

OpenStudy (anonymous):

1 2 3 0 those are the choices

OpenStudy (sirm3d):

1.

OpenStudy (anonymous):

it looks just like \(y=x^3\) only shifted to the left 8 units and up 9 units \(y=x^3\) only has one zero we can find it if you like

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

\[(x+8)^3+9=0\] \[(x+8)^3=-9\] \[x+8=\sqrt[3]{-9}\] \[x=-8+\sqrt[3]{-9}\]

OpenStudy (anonymous):

thanks:)